OpenStudy (anonymous):

The remainder when 7^89761 is divided by 15 is?

OpenStudy (anonymous):

@experimentX @phi @satellite73 @saifoo.khan @apoorvk @yash2651995

OpenStudy (saifoo.khan):

calc?

OpenStudy (anonymous):

Nop........there should be some otherway! calculating is impossible

OpenStudy (saifoo.khan):

So we have to solve with calculation?

OpenStudy (experimentx):

try to find some cylic pattern

OpenStudy (lgbasallote):

must have something to do with the exponent

OpenStudy (anonymous):

@satellite73 i got the pattern but the answer should be 7 there it has written1!!

OpenStudy (anonymous):

@phi

OpenStudy (experimentx):

lol ... i remember that yes i guess it should be 7

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=7%5E89761+mod+15

OpenStudy (btaylor):

There is a trend when dividing powers of 7 by 15. 7^0/15 = R1 7^1/15 = R7 7^2/15 = R4 7^3/15 = R13 7^4/15 = R1 7^5/15 = R7 7^6/15 = R4 7^7/15 = R13 So 7^(4n) will have a remainder of 1, 7^(4n+1) will have a remainder 7, and so on. 89761 is a 4n+1 type.

OpenStudy (anonymous):

@experimentX how 7 i too got 1 plz help

OpenStudy (experimentx):

7^(89761 mod 4) mod 15 89761 mod 4 = 1 so, 7^1 mod 15 = 7

OpenStudy (anonymous):

@experimentX is that mod is absolute value

OpenStudy (experimentx):

no ... modulus remainder when you divide some number by some number

OpenStudy (experimentx):

try experimenting http://www.wolframalpha.com/input/?i=7+mod+5

OpenStudy (anonymous):

can u detaily write it becoz it is an imp step i am confused

OpenStudy (experimentx):

Oh sorry ... it's called modulo ,,, or so

OpenStudy (anonymous):